OpenStudy (anonymous):

Find the average rate of change of each function on the interval specified. Your answers will be expressions. j(x) = 3x^3 on [1, 1+h]

OpenStudy (anonymous):

compute \[\frac{j(1+h)-j(1)}{h}\]

OpenStudy (anonymous):

average rate of change is defined as y2-y1/x2-x1 right?

OpenStudy (anonymous):

would it be j(1) = 3(1)^3 and j(1+h) = 3(1+h)^3

OpenStudy (anonymous):

so what I got was j(1) = 3(1)^3 = 3 j(1+h)^3 = 3h^3+3h^2+3h+1=3(h^3+3h^2+3h+1)=3h^3+9h^2+9h+3